Exposure to benzene emissions from automobile traffic is known to increase the risk of leukaemia in adults. It is not known whether children have a similar risk from this exposure. A cohort study was initiated to examine the effect of exposure to benzene emissions from automobile traffic on the risk of leukaemia in people aged 0 to 14 years. There were 316people in the cohort exposed to benzene emissions from automobile traffic. There were 96people in the cohort not exposed to benzene emissions.

The number of cases of leukaemia in the two cohorts was compared after ten years. There were 64 cases in the exposed cohort and 16 cases in the unexposed cohort.

  1. (a) What is the rate of disease in these two groups? (three marks)

  2. (b) Identify the correct measure of effect between exposure and outcome for this study? (two

    mark)

  3. (c) What is the numerical value of this measure? (two marks)

  4. (d) Interpret this result in one sentence using plain language. (three marks)
Answer #1

We can form the table as follows from the information given:

Disease No Disease Total
Exposed 64 252 316
Non Exposed 16 80 96
Total 80 332 412

a) Rate of disease in the exposed group = 64/316 = 0.20

Rate of disease in the not exposed group = 16/96 = 0.167

b) The correct measure of effect is Risk Ratio

c) Risk Ratio is calculated by:

RR = 0.20/0.167 = 1.20

d) Interpretation: Leukaemia in the exposed group is 1.20 times more likely to occur than in the non-exposed group